OpenStudy (anonymous):

the sum of the page numbers on consecutive pages of a book is 267. What are the page numbers?

OpenStudy (asnaseer):

lets call one of the pages 'p'. then the next consecutive one must be 'p+1'. you are given their sum, so we know:\[p+p+1=267\]you should be able to solve from there...

OpenStudy (anonymous):

i need more help....

OpenStudy (asnaseer):

which part are you stuck on?

OpenStudy (anonymous):

how am i suppose to find what the P are.?

OpenStudy (anonymous):

am i suppose to divide?

OpenStudy (asnaseer):

ok, lets break it down into smaller parts. first, how can we simplify: p + p

OpenStudy (asnaseer):

how many p's do we end up with?

OpenStudy (anonymous):

2 p's?!

OpenStudy (asnaseer):

correct, so we can now write the equation as: 2p + 1 = 267 next, take '1' away from both sides

OpenStudy (anonymous):

2p+1=267 1 1 ----------- 1p=166?

OpenStudy (asnaseer):

no - that is not the right way of subtracting. we do it like this: 2p + 1 - 1 = 267 - 1 to get: 2p = 266

OpenStudy (asnaseer):

next we divide both sides by 2

OpenStudy (anonymous):

which equals 133?

OpenStudy (anonymous):

so 1p = 133?!!!

OpenStudy (asnaseer):

correct - so we end up with: p = 133 so now we know the page numbers are 133 and 134

OpenStudy (anonymous):

yay!!!! so 133 and 134 is my answer!?

OpenStudy (asnaseer):

correct